# Thin gap approximations for microfluidic device design

### v1

@@ 2 A new perspective of the Hele-Shaw approximation
Our closure method follows the general framework of MWR Finlayson (2013). In the standard application of MWR to obtain a reduced-order model for a differential equation $\mathcal{L}(u)=f$, one begins by selecting a set of basis functions ${\phi_{i}(z)}_{i=1}^{N}$ and approximating the solution as $u(x,z)\approx\sum_{i=1}^{N}u_{i}(y)\phi_{i}(z)$. The residual is then defined as $r(y,z)=\mathcal{L}(u(y,z))-f(y,z)$. A set of weight functions ${w_{j}(z)}_{j=1}^{N}$ is chosen, and the residual is projected onto these weights via $\int_{\Omega}w_{j}(z)r(y,z)\mathrm{d}z=0$, $j=1,\dots,N$, yielding a system of $N$ equations for the unknown coefficients $u_{i}(y)$}. Our modification lies in the third step: rather than integrating the equation against multiple orthogonal weights, we assume a closure equation with unknown parameters, substitute it into the residual, and determine the parameters by minimizing the weighted $L^{2}$ norm of the resulting expression. This approach offers greater flexibility in the form of the closure equation.
Ideally, if minimizing the residual directly yielded an optimal approximation to the true solution—without requiring the original equation to be solved—this would be a highly desirable outcome. However, this is not generally the case. In our example (Fig. 2), the near-optimal approximation results from a carefully chosen combination of basis functions and residual metric that reflect the underlying physical behavior of the system. We use Gegenbauer polynomials as basis functions, motivated by the parabolic profile of the classical Hele-Shaw approximation. The weighted $L^{2}$ norm used for residual minimization is chosen to align with the orthogonality properties of these polynomials.
@@
Stokes flow in coaxial flow devices: Under the assumption of Stokes flow, the nonlinear term $\mathbf{u}\cdot\nabla\mathbf{u}$ is again negligible. The optimal 2D approximation replaces the operator $\Delta$ by $a\Delta_{H}-12$, where $\Delta_{H}\equiv\partial_{x}^{2}+\partial_{y}^{2}$, even when the flow is three dimensional. The same choice of $a$ may be justified (residuals such Eq. (2) are obtained for each component of the fully 3D Stokes flow). The residual analysis of the continuity equation yields $\partial_{x}u_{1}+\partial_{y}v_{1}=0$. We test the approximation on a thin-gap but non-rectilinear flow that is commonly realized in microfluidic devices. In coaxial flow, three fluid streams, flowing sufficiently slowly that $\mathrm{\mathrm{Re}}=0$, are combined and pass through a narrow aperture (Fig. 3a, Anna et al. (2003)), emerging as a inner fluid thread sandwiched between the outer fluids. In real coaxial flow devices, the surface tension of the inner fluid thread breaks it up into droplets. We consider a version of the coaxial flow experiment without viscosity contrasts or surface tension, focusing only on the ability of the optimal 2D model to render the geometry of the interface between the inner and outer fluids. We used COMSOL Multiphysics 5.4 to solve both the 3D Stokes equations and the 2D approximate models, and to compare their streamlines with those from the optimal 2D approximation.
Simulations were performed using the geometry from Anna et al. (2003), while varying the mass flux ratio $Q$ between the inner and outer flows. At a representative value of $Q$, (Fig. 3b), the optimal 2D approximation with $a=6/5$ closely approximates the interface in both the narrowing and diverging sections of the channel, whereas the classical Hele-Shaw approximation ($a=0$) overpredicts the downstream width of the middle layer, and provides only qualitative agreement in shape.
@@
Figure 3: Coaxial flow. (a) Simulation geometry based on Anna et al. (2003): left— inlets, right- outlet. Outer inlets have mass flux $Q$, the inner inlet has flux 11. Inlet and outlet lengths are extended by 20 to ensure fully developed flow. (b) Dividing streamlines for $Q=1.7$: classical Hele–Shaw (red), derived $a=6/5$ approximation (blue dotted), and 3D simulation (black dashed).
@@
The agreement between the 2D approximation and the full 3D computation in the far downstream region is expected for two reasons. First, the flow becomes rectilinear, resembling Poiseuille flow—a regime in which our optimal 2D approximation is already known to be accurate (Fig. 2b). Second, mass conservation between the inlets and the outlet constrains the position of the interfaces within the downstream Poiseuille flow. Indeed, it is unsurprising that the classical Hele-Shaw approximation will over-predict the width of the inner-jet (Fig. 3b), because it treats the downstream flow as uniform or plug-like in $y$, omitting the slowdown near the sidewalls of the channel that prevents narrowing of the outer flows. However, the optimal 2D approximation renders the interface accurately even where the flow passes through the narrow orifice separating the inlets and outlet. Although the orifice is relatively narrow (aspect ratio, width:depth <0.4<0.4), and conventional wisdom holds that Hele-Shaw-type models are accurate only when the channel is much wider (aspect ratio $\gtrsim$ 1), the optimal 2D approximation still performs well. Remarkably, the flow remains approximately two-dimensional even within the orifice: the ratio of out-of-plane to in-plane velocity components peaks at just 0.22. This modest degree of three-dimensionality, forced by the thin gap geometry, suggests that the validity of our optimal 2D model extends well beyond the classical limits typically associated with the Hele-Shaw approximation.
@@ 3 Thin gap flows at moderate Reynolds numbers
After validating the accuracy of the approximation for Stokes flow (i.e. for thin gap microfluidic devices), we can generalize it to the case of thin gap flows at moderate Reynolds numbers. These conditions are often realized in inertial microfluidic devices, which are commonly built using the same soft-lithography techniques, and similar dimensions to regular microfluidic devices, but are perfused with high pressure pumps, to create $\sim$ m/s flows Di Carlo (2009).
@@
$\small\mathrm{Re}(a\partial_{t}\mathbf{u}_{H}+b\mathbf{u}_{H}\cdot\nabla_{H}\mathbf{u}_{H})=c\Delta_{H}\mathbf{u}_{H}-d\mathbf{u}_{H}-\nabla_{H}p,\;\nabla_{H}\cdot\mathbf{u}_{H}=0,\;\left.\mathbf{u}_{H}\right|_{\partial\Omega}=0.$ (6)
Following the same weighted residual analysis, using the weight functions $w(z)=\frac{1}{4}-z^{2}$ and $w(z)=1$ yields the following sets of coefficients, respectively:
@@
Previous work has reported these and similar closure models, arrived at by different approximations and techniques. For example, Plouraboué & Hinch (2002); Yuan & Azaiez (2014) used $a=1$, $b=1$, $c=1$, and $d=12$, while for the case of Stokes flow, Zeng et al. (2003) proposed $a=b=0$, $c=\frac{12}{\pi^{2}}\approx 1.21$, and $d=12$. The closure equation with $w(z)=1$ coefficients was derived in Gondret & Rabaud (1997) by assuming a parabolic flow profile, and averaging the Navier-Stokes equations across the gap.
@@
By positing an expansion of the velocity fields as a (non-spectral) sequence of polynomials, and matching powers of $z$, Ruyer-Quil (2001) derived a version of the closure equation with values $a=\frac{6}{5}$, $b=\frac{54}{35}$, $c=0$, and $d=12$. Although it omits in-plane viscous stresses, this model has found wide use in studies of fluid instability in Hele-Shaw geometries Chevalier et al. (2006). Martin et al. (2002) add in-plane viscous stresses with an effective viscosity to this model resulting in identical coefficients to Eq. (7a). More recently, Li et al. (2021) arrived at the same 2D approximation as (7a), via a lubrication approximation approach, viz. systematic expansion of all pressure and velocity fields in powers of $h/L$. By calculating terms up to $O(h^{2}/L^{2})$, and neglecting terms proportional to the square of the Reynolds number, they arrive at Eq. (7a). We do not therefore make any claim to be the first to arrive at this 2D approximation, but note that optimal closure via MWR provides a direct and physically transparent route to deriving the equation.
@@
Figure 4: 2D approximation for finite $\mathrm{Re}$ flows in a centrifuge-on-a-chip geometry. Comparison of the separation bubble width in 3D simulation (black), 2D approximation (blue, Eq. (7a)), and unweighted residual approximation (red, Eq. (7b)). Right panels: Streamline patterns colored by velocity magnitude (red = larger, blue = smaller), comparing the 2D approximation (top) with the full 3D simulation (bottom).
@@
To test our 2D approximation and to compare weighted and unweighted residuals, we compare both 2D approximations with 3D simulations of flow in a centrifuge-on-a-chip (Fig. 1c), which consists of a rectangular chamber connected to two narrow, elongated channels serving as the inlet and outlet Hur et al. (2011). In our simulations, performed in COMSOL Multiphysics 6.3, the chamber occupies $(x,y,z)\in[-50,50]\times[-50,50]\times[-0.5,0.5]$, while the inlet and outlet have unit square cross-sections. The inlet region of the flow resembles the channel expansion on which Li et al. (2021) have already validated the 2D approximation (7a), but the centrifuge-on-a-chip is a paradigmatic example of the design of inertial microfluidic devices to fractionate particles, such as cells in a blood sample, by size, so centrifuge-on-a-chip flows merit separate attention.
@@
As the Reynolds number $\mathrm{Re}=Uh/\nu$ of the inlet flow is increased, a separation eddy forms at the centrifuge entrance, and this eddy plays a role in trapping of large particles Paiè et al. (2017). We measure the eddy size by interpolating the $v=0$ contour to the wall x=50x=-50. The 2D approximation (7a) (left panel, blue diamonds) models the growth of the 3D eddy with $\mathrm{Re}$ (black circles), better than the unweighted residual model (red circles) (7b). Though both approximations slightly overpredict the value of $\mathrm{Re}$ at which separation first occurs (17.5 compared to 16 for the fully 3D simulation), the unweighted residual approximation underestimates the separation eddy size by a factor of 16-21% over the entire $\mathrm{Re}$-range (Fig. 4), while the error from our 2D approximation (7a) never exceeds 12%. Comparing flow fields (Fig.4, right panel), we see that the 2D approximation from (7a) also quantitatively matches the streamlines and velocities obtained in 3D simulations.
Figure 5: Fidelity of the numerically computed 3D flow field to the Hele-Shaw (parabolic flow profile) for a centrifuge-on-a-chip flow. (a) $r$ (defined in Eq. (8)) at $\mathrm{Re}=200$ across the centrifuge chamber. (b,c) $u$- profiles against $z$ sampled on two different transects (locations shown in (a)). To compare shapes, all profiles are normalized so that $u=1$ at $z=0$ (b) shows a small $r$-region, where the Hele-Shaw approximation is valid, and (c) a region on the boundary of the inlet jet, where it is not. Line colors give $x$-ordinates. (d) Adding $Q_{3}$ to our basis allows non-Hele-Shaw features to be rendered. Two specific non-parabolic velocity profiles from the $y=6$ transect are shown (solid lines) along with their projections onto $Q_{1}$ and $Q_{3}$ (dotted line). Simulations were performed in COMSOL Multiphysics 6.2, with P2+P2 elements and the extra fine mesh further refined to ensure a maximum cell length of 0.1.
4 Breakdown of the 2D model

### v2

@@ 2.2 Method of weighted residuals
In the standard application of MWR to obtain a reduced-order model for a differential equation $\mathcal{L}(u)=f$, one first (i) selects a set of basis functions $\left\{\phi_{n}(z)\right\}_{n=1}^{N}$ and approximates the solution as $u(x,z)\approx\sum_{n=1}^{N}u_{n}(y)\phi_{n}(z)$. Next (ii), the residual is then defined as $r(y,z)=\mathcal{L}(u(y,z))-f(y,z)$. Finally (iii), a set of weight functions $\left\{w_{n}(z)\right\}_{n=1}^{N}$ is chosen, and the residual is projected onto these weights via $\int_{\Omega}w_{n}(z)r(y,z)\mathrm{d}z=0$, $j=1,\dots,N$, yielding a system of $N$ equations for the unknown coefficients $u_{n}(y)$}.
We can express both approximations derived in the previous section within the MWR framework. First, we take the weight functions to be $w_{n}=C_{n}^{(1/2)}(2z)$. Owing to the orthogonality of the Gegenbauer polynomials, imposing $\int_{\Omega}w_{n}(z)r(y,z)\mathrm{d}z=0$ is equivalent to expanding the residual in the $C_{n}^{(1/2)}$ basis and setting the corresponding coefficients to zero. Expanding the residual in (4) in terms of $C_{n}^{(1/2)}$ yields
$r(y,z)=\left(\partial_{y}^{2}u_{0}(y)-12u_{0}(y)+12\right)C_{0}^{(1/2)}(2z)-\partial_{y}^{2}u_{0}(y)C_{2}^{(1/2)}(2z),$ (16)
[…]
The coefficient of $C_{0}^{(1/2)}$ recovers the approximation (8) with $a=1$.
Next, we take $w_{n}=(1-4z^{2})C_{n}^{(3/2)}(2z)$ as the weight functions. Expanding the same residual in the $C_{n}^{(3/2)}$ basis yields
$r(y,z)=\left(\tfrac{6}{5}\partial_{y}^{2}u_{0}(y)-12u_{0}(y)+12\right)C_{0}^{(3/2)}(2z)-\tfrac{1}{5}\partial_{y}^{2}u_{0}(y)C_{2}^{(3/2)}(2z),$ (17)
[…]
The coefficient of $C_{0}^{(3/2)}$ corresponds exactly to the approximation (10) with $a=6/5$.
In this sense, the derivation in the previous section is fully consistent with the MWR formulation: the velocity field is expanded in $C_{n}^{(-1/2)}(2z)$, while the residual is projected onto the space spanned by $C_{n}^{(3/2)}(2z)$. The resulting procedure may therefore be interpreted as a Petrov–Galerkin method, since the trial and test spaces are constructed from distinct polynomial families.
Although the optimal closure equation can be obtained within the classical MWR framework by selecting appropriate basis functions and weights, our discussion in the previous section highlights two nontrivial points. First, we demonstrate an alternative to the third step in MWR: instead of integrating the residual against multiple orthogonal weights, we propose a closure equation with unknown parameters, substitute it into the residual, and determine the parameters by minimizing the weighted $L^{2}$ norm. This formulation provides greater flexibility in the closure form and may lead to different relations in other contexts. Second, owing to the completeness of the basis, the systems obtained by projecting the residual onto Legendre polynomials $C_{n}^{(1/2)}$, Gegenbauer polynomials $C_{n}^{(3/2)}$, or even sine functions (as in Pepper et al. (2010)), become asymptotically equivalent when many terms are retained in the velocity expansion. However, when only a few terms are used, the choice of weight strongly influences the approximation accuracy. In such cases, one must carefully select both the basis functions and the residual metric to capture the underlying physical behavior of the system.
2.3 Stokes flow in coaxial flow devices
Under the assumption of Stokes flow, the nonlinear term $\mathbf{u}\cdot\nabla\mathbf{u}$ is again negligible. The optimal 2D approximation replaces the operator $\Delta$ by $a\Delta_{H}-12$, where $\Delta_{H}\equiv\partial_{x}^{2}+\partial_{y}^{2}$, even when the flow is three dimensional. The same choice of $a$ may be justified (residuals resembling Eq. (7) are obtained for each component of the fully 3D Stokes flow). The residual analysis of the continuity equation yields $\partial_{x}u_{0}+\partial_{y}v_{0}=0$.
@@ 2.3 Stokes flow in coaxial flow devices
We test the approximation on a thin-gap but non-rectilinear flow that is commonly realized in microfluidic devices. In coaxial flow, three fluid streams, flowing sufficiently slowly that $\mathrm{\mathrm{Re}}=0$, are combined and pass through a narrow aperture (Fig. 3a, Anna et al. (2003)), emerging as a inner fluid thread sandwiched between the outer fluids. In real coaxial flow devices, the surface tension of the inner fluid thread breaks it up into droplets. We consider a version of the coaxial flow experiment without viscosity contrasts or surface tension, focusing only on the ability of the optimal 2D model to render the geometry of the interface between the inner and outer fluids. We used COMSOL Multiphysics 5.4 to solve both the 3D Stokes equations and the 2D approximate models, and to compare their streamlines with those from the optimal 2D approximation. Simulations were performed using the geometry from Anna et al. (2003), while varying the mass flux ratio $Q$ between the inner and outer flows. At a representative value of $Q$, (Fig. 3b), the optimal 2D approximation with $a=6/5$ closely approximates the interface in both the narrowing and diverging sections of the channel, whereas the classical Hele-Shaw approximation ($a=0$) overpredicts the downstream width of the middle layer, and provides only qualitative agreement in shape.
@@
Figure 3: Coaxial flow. (a) Simulation geometry based on Anna et al. (2003): left: inlets, right: outlet. The outer inlets have mass flux $Q$, the inner inlet has flux 11. Inlet and outlet lengths are extended by 20 to ensure fully developed flow. (b) Dividing streamlines for $Q=1.7$: classical Hele–Shaw (red dashed), derived $a=6/5$ approximation (blue dotted), and 3D simulation (black dashed).
@@
The agreement between the 2D approximation and the full 3D computation in the far downstream region is expected for two reasons. First, the flow becomes rectilinear, resembling Poiseuille flow—a regime in which our optimal 2D approximation is already known to be accurate (Fig. 2b). Second, mass conservation between the inlets and the outlet constrains the position of the interfaces within the downstream Poiseuille flow. Indeed, it is unsurprising that the classical Hele-Shaw approximation will over-predict the width of the inner-jet (Fig. 3b), because it treats the downstream flow as uniform or plug-like in $y$, omitting the slowdown near the sidewalls of the channel that prevents narrowing of the outer flows. However, the optimal 2D approximation renders the interface accurately even through the narrow orifice separating the inlets and outlet. Although the orifice is relatively narrow (aspect ratio, width: depth <0.4<0.4), and conventional wisdom holds that Hele-Shaw-type models are accurate only when the channel is much wider (aspect ratio $\gtrsim$ 1), the optimal 2D approximation still performs well. Remarkably, the flow remains approximately two-dimensional even within the orifice: the ratio of out-of-plane to in-plane velocity components peaks at just 0.22. This modest degree of three-dimensionality, forced by the thin gap geometry, suggests that the validity of our optimal 2D model extends well beyond the classical limits typically associated with the Hele-Shaw approximation.
@@ 3 Thin gap flows at moderate Reynolds numbers
After validating the accuracy of the approximation for Stokes flow (i.e. for thin gap microfluidic devices), we can generalize it to the case of thin gap flows at moderate Reynolds numbers. These conditions are often realized in inertial microfluidic devices, which can be built using the same soft-lithography techniques, and similar dimensions to regular microfluidic devices, but are perfused with high pressure pumps, to create $\sim$ m/s flows (see Di Carlo (2009)).
@@
$\small\mathrm{Re}(a\partial_{t}\mathbf{u}_{H}+b\mathbf{u}_{H}\cdot\nabla_{H}\mathbf{u}_{H})=c\Delta_{H}\mathbf{u}_{H}-\frac{d}{H^{2}}\mathbf{u}_{H}-\nabla_{H}p,\;\nabla_{H}\cdot\mathbf{u}_{H}=0,\;\left.\mathbf{u}_{H}\right|_{\partial\Omega}=0.$ (18)
Here, $\mathrm{Re}=\tfrac{LU}{\nu}$ is the Reynolds number, with $L$ the characteristic length scale, $U$ the characteristic velocity, and $\nu$ the kinematic viscosity. The choice of $L$ influences the nondimensionalization. For example, taking $L=\tfrac{h}{2}$ (half the channel depth) yields the vertical domain $z\in[-1,1]$, while taking $L=h$ (the full depth, as in this work) rescales the domain to $z\in[-\tfrac{1}{2},\tfrac{1}{2}]$. To accommodate both conventions, we denote by $H$ the nondimensionalized gap width.
Following the same weighted residual analysis, using the weight functions $w(z)=1-4z^{2}$ and $w(z)=1$ yields, respectively:
@@
Previous work has reported these and similar closure models, arrived at by different approximations and techniques. For example, Plouraboué & Hinch (2002); Yuan & Azaiez (2014) used $a=1$, $b=1$, $c=1$, and $d=12$, while for the case of Stokes flow, Zeng et al. (2003) proposed $a=b=0$, $c=\frac{12}{\pi^{2}}\approx 1.21$, and $d=12$. The closure equation with $w(z)=1$ coefficients in Eq (19b) was derived in Gondret & Rabaud (1997) by assuming a parabolic flow profile, and averaging the Navier-Stokes equations across the gap.
@@
Ruyer-Quil (2001) first demonstrated that inertial effects give rise to the coefficient $\tfrac{54}{35}$ in the depth-averaged equation obtained via the MWR. The present work differs in two important respects. First, the coefficient in Ruyer-Quil (2001) was derived by expanding the velocity field as a sequence of polynomials of the form $z^{j}(1-4z^{2})$ and matching coefficients of like powers of $z$. This non-orthogonal polynomial basis can hinder the derivation of higher-order approximations, leading to more strongly coupled systems with reduced accuracy. Second, the approximation in Ruyer-Quil (2001) neglects in-plane viscous stresses. Despite this limitation, the model has been widely adopted in studies of fluid instabilities in Hele–Shaw geometries (Chevalier et al. (2006)). Martin et al. (2002) add in-plane viscous stresses with an effective viscosity resulting in identical coefficients to Eq. (19a). More recently, Li et al. (2021) performed an asymptotic analysis, expanding the pressure and velocity fields in powers of $h/L$. By retaining terms up to $O(h^{2}/L^{2})$, discarding terms proportional to the square of the Reynolds number, and several additional approximation assumptions, they obtained the same 2D approximation as (19a). We do not, therefore, make any claim to be the first to arrive at this 2D approximation, but note that optimal closure via MWR provides a direct and physically transparent route to deriving the equation, and its higher order approximations.
@@
Figure 4: 2D approximation for finite $\mathrm{Re}$ flows in a centrifuge-on-a-chip geometry. Comparison of the separation eddy size in 3D simulation (black), 2D approximation (blue, Eq. (19a)), second order 2D approximation (orange, Eq. (20)), and unweighted residual approximation (red, Eq. (19b)). Right panels: Streamline patterns colored by velocity magnitude (red = larger, blue = smaller), comparing the 2D approximation (top) with the full 3D simulation (bottom).
@@
To compare weighted and unweighted residuals, we compare both 2D approximations with 3D simulations of flow in a centrifuge-on-a-chip (Fig. 1c, Hur et al. (2011)), which consists of a rectangular chamber connected to two narrow, elongated channels serving as the inlet and outlet. In our simulations, performed in COMSOL Multiphysics 6.3, the chamber occupies $(x,y,z)\in[-50,50]\times[-50,50]\times[-0.5,0.5]$, while the inlet and outlet have unit square cross-sections. The inlet region of the flow resembles the channel expansion on which Li et al. (2021) has already validated the 2D approximation (19a), but the centrifuge-on-a-chip is a paradigmatic example of the design of inertial microfluidic devices to fractionate particles, such as cells in a blood sample, by size, so centrifuge-on-a-chip flows merit separate attention.
@@
As the Reynolds number $\mathrm{Re}=Uh/\nu$ of the inlet flow is increased, a separation eddy forms at the centrifuge entrance, and this eddy plays a role in trapping of large particles (Paiè et al. (2017)). We measure the eddy size by interpolating the $v=0$ contour to the wall x=50x=-50. The 2D approximation (19a) (left panel, blue diamonds) models the growth of the 3D eddy with $\mathrm{Re}$ (black circles), better than the unweighted residual model (red circles) (19b). Though both approximations slightly overpredict the value of $\mathrm{Re}$ at which separation first occurs (17.5 compared to 16 for the fully 3D simulation), the unweighted residual approximation underestimates the separation eddy size by a factor of 16-21% over the entire $\mathrm{Re}$-range (Fig. 4), while the error from our 2D approximation (19a) never exceeds 12%. Comparing flow fields (Fig.4, right panel), we see that the 2D approximation from (19a) also quantitatively matches the streamlines and velocities obtained in 3D simulations.
4 Higher order corrections